OpenStudy (anonymous):

PLEASE HELP One year there was a total of 42 commercial and noncommercial orbital launches worldwide. In addition the number of noncommercial launches was three more than twice the number of commercial orbital launches. Determine the number of commercial and noncommercial orbital launches.

OpenStudy (anonymous):

answer is 13 commercial and 30 noncommercial

OpenStudy (anonymous):

your question is not clear

OpenStudy (anonymous):

please explain the solution. @flutterflies

OpenStudy (anonymous):

n=2c+3 c+(2c+3)=42 3c+3 = 42 3c= 39 c= 13 plug in for c 13+n=42 n=30 13 commercial and 30 noncommercials

OpenStudy (anonymous):

@gk_goel

OpenStudy (anonymous):

n=29.....13commercial and 29 non-commercial..... @flutterflies

OpenStudy (anonymous):

opps thanks @gk_goel